# The Energy Cost of Reasoning: Analyzing Energy Usage in LLMs with Test-time Compute

### v1

@@ Abstract
Scaling large language models (LLMs) has driven significant advancements, yet it faces diminishing returns and escalating energy demands. This work introduces test-time compute (TTC)—allocating additional computational resources during inference—as a compelling complement to conventional scaling strategies. Specifically, we investigate whether employing TTC can achieve superior accuracy-energy trade-offs compared to simply increasing model size. Our empirical analysis reveals that TTC surpasses traditional model scaling in accuracy/energy efficiency, with notable gains in tasks demanding complex reasoning rather than mere factual recall. Further, we identify a critical interaction between TTC performance and output sequence length, demonstrating that strategically adjusting compute resources at inference time according to query complexity can substantially enhance efficiency. Our findings advocate for TTC as a promising direction, enabling more sustainable, accurate, and adaptable deployment of future language models without incurring additional pretraining costs.
@@ 1 Introduction
Test-time compute (TTC) has emerged as a promising approach to significantly enhance the performance of large language models (LLMs) without the reliance on traditional scaling laws. Conventionally, improving neural language models involved scaling either the number of parameters or expanding the training dataset [2020arXiv200108361K]. However, this approach demands substantial computational resources, resulting in considerable energy consumption during both training and inference phases. Furthermore, such scaling is increasingly constrained by the limited availability of sufficiently large and diverse datasets[2022arXiv221104325V], highlighting diminishing returns in terms of resource efficiency.
@@
Recent research addresses this limitation by proposing a new strategy to enhance the performance of LLMs: allocating additional computational resources at inference time, known as TTC. TTC is inspired by human reasoning process, where humans reflect and reason logically to answer challenging questions rather than responding impulsively. Similarly, LLMs should leverage additional computational effort during inference to improve their responses. Prominent examples of TTC methodologies include prompting techniques like Chain-of-Thought (CoT) reasoning  [2022arXiv220311171W, 2023arXiv230510601Y], self-revision methods [2024arXiv240601297K, 2024arXiv240615673L, 2023arXiv230317651M], post-training for reasoning [deepseek, star, rest], and ensemble-like approaches aggregating response from multiple parallel generations to improve the final response [llmonk, getting_50, 2023arXiv230306135I, 2022arXiv220311171W]. Despite methodological differences, these strategies collectively emphasize additional computational investment during inference to enhance model accuracy.
@@
Nevertheless, the rising computational demands of TTC necessitate careful consideration of associated energy costs. Recent industry reports underscore this concern, with Meta attributing up to 70% of its AI power consumption to inference processes [wu2022sustainable], Google reporting 60% of machine learning energy usage [patterson2022carbon], and AWS indicating inference-related demands accounting for 80-90% of computing resources [barr2019amazon] all without TTC. Moreover, broader data center energy consumption trends indicate a rapid escalation, expected to constitute up to 12% of total U.S. electricity usage by 2028 [shehabi2024united]. The environmental impact and financial implications of such intensive energy consumption underline the urgency of evaluating the sustainability and efficiency of TTC approaches.
@@
In this paper, we critically examine the performance-energy trade-offs associated with TTC compared to traditional model scaling. Specifically, we define performance as accuracy of evaluating benchmarks and investigate the relationship between accuracy improvements and energy consumption incurred by employing TTC methods versus scaling model sizes. Through comprehensive experiments conducted on Qwen2.5 [qwen2.5] models (1.5B, 7B, 14B, and 32B parameters) using A100 GPUs across mathematical, coding, and common-sense reasoning benchmarks, we derive key insights into the comparative advantages and costs of TTC.
@@
  • TTC demonstrates superior accuracy-energy trade-off improvements over traditional model scaling especially for performant models. Notably, smaller models enhanced with TTC can outperform substantially larger models relying solely on scale.
  • However, indiscriminate application of TTC can drastically elevate energy consumption, with cases where TTC usage amplifies energy costs up to 97.48$\times$, as observed with the 7B model on a coding benchmark.
@@
  • Complex questions inherently drive models to allocate greater computational resources during inference, paralleling human cognitive processes in handling challenging tasks.
To the best of our knowledge, this study represents one of the first systematic explorations into the energy implications of TTC for LLM inference. Our analysis provides a foundational understanding of how TTC can serve as a more energy-efficient alternative and complement to traditional scaling while also underscoring the potential for substantial resource expenditure. We explicitly focus on inference-related impacts, leaving exploration of training-phase considerations as important avenues for future research.
Figure 1: Accuracy vs Energy per query. Each dot in the line represent Qwen2.5 1.5B, 7B, 14B, and 32B from left to right, respectively.
@@ 2 Background
Inference for transformer-based LLMs [llama, qwen2.5, gemini, gpt4] typically involves two sequential stages: prefill and decode. Initially, when an input sequence is provided, it undergoes the prefill stage, where all tokens in the seqeunce are processed simultaneously. This stage computes the first output token and generates a Key-Value cache (KV-cache), which stores contextual information from the input sequence. Following the prefill stage, the model enters the decode stage, wherein it autoregressively generates subsequent tokens one-by-one, leveraging the previously generated KV-cache.
Each of these stages presents distinct computational bottlenecks. The prefill stage is predominantly compute-bound due to extensive matrix multiplications resulting from processing all tokens simultaneously. In contrast, the decode stage typically becomes memory bandwidth (BW)-bound because of the frequent accesses to the KV-cache. Recognizing this distinction, state-of-the-art LLM inference engines optimize latency and throughput by overlapping these two stages [orca, vllm, sglang, nanoflow, deepspeed], efficiently exploiting their differing computational characteristics within the same model architecture.
@@
Recently, leveraging additional computational resources at inference time—known as TTC—has emerged as an effective strategy for enhancing the reasoning capabilities and overall accuracy of LLMs. We categorize TTC techniques into two categories based on their resource utilization patterns: 1) techniques that increase the input tokens more [llmonk, 2022arXiv220311171W, few_shot] or 2) those that increase output tokens more [deepseek, 2023arXiv230317651M, star, rest]. Each category uniquely affects the resource demands of the inference pipeline.
Increasing the number of input tokens primarily impacts the prefill stage, significantly intensifying its computational load, as the compute requirement scales quadratically with the length of the input sequence. On the other hand, increasing the number of output tokens predominantly affects the decode stage. Specifically, extending the output length linearly increases both the size of the KV-cache and the number of autoregressive decoding steps, thus elevating memory BW requirements.
@@
While considerable research has focused on improving the performance and reasoning capabilities of LLMs, the energy consumption associated with their deployment and usage remains relatively underexplored. Typically, energy consumption is measured by tracking power usage throughout model runtime and integrating this measurement over time.
Early pioneering efforts, such as the study by Strubell et al. [strubell2020energy], began to systematically investigate the energy cost of training transformer models, spurring subsequent studies that quantified energy consumption during model training [patterson2022carbon]. With the increasing deployment of LLMs in practical applications, recent studies have expanded this scope to include energy measurements specific to inference workloads [luccioni2024power, fernandez2025energy, wu2025unveiling, patel2024characterizing, stojkovic2024towards].
@@ 3 Methodology
Table 1: List of benchmarks used in this study
Task Benchmark Task Benchmark Task Benchmark
Math AIME24 [aime24] Code HumanEval [humaneval] Common Sense HellaSwag [hellaswag]
AIME25 [aime25] LiveCodeBench [livecodebench] MMLU [mmlu]
GPQADiamond [gpqa] MBPP [mbpp] GSM8K [gsm8k]
Math500 [math500] CodeForces [codeforces] CommonsenseQA [talmor-etal-2019-commonsenseqa]
In this study, we systematically measure the additional energy consumption associated with TTC methods in the context of LLM inference. To quantify this energy overhead, we base our work on Evalchemy [evalchemy] and utilize SGLang [sglang] for model execution on NVIDIA A100 GPUs [a100], each with a maximum power rating of 500W. We run each benchmark from start to end assuming that all questions are present to be processed to eliminate any scheduling overhead or queuing delay. Power consumption during model inference is monitored using the NVIDIA Management Library (NVML) [nvml]. We integrate power usage over the inference period to compute total energy consumption for each model evaluation.
@@
We group TTC approaches into two distinct categories: (1) methods that increase the number of input tokens and (2) methods that expand the number of output tokens. For the first category, we select parallel sampling [llmonk, getting_50, 2023arXiv230306135I, 2022arXiv220311171W] with majority vote (MV) to finalize the answer. Although parallel sampling can leverage prefix caching by performing the prefill stage only once, we deliberately opt to repeat the prefill stage for each sample independently. This strategy ensures effective utilization of varying sampling temperatures, thereby producing a diverse set of candidate outputs.
For the second category, we adopt the reasoning token (RT) approach [deepseek, star, rest]. Unlike baseline models that lack post-training enhancements and thus cannot effectively refine their outputs [2023arXiv231001798H], models utilizing reasoning tokens have demonstrated a notable capacity for self-correction and improved reasoning. This enhanced capability results from specialized tokens learned during the post-training phase.
@@
We base our study on models supported by the DeepSeek-R1 framework [deepseek], selecting four representative models from the Qwen2.5 family [qwen2.5]: 1.5B, 7B, 14B, and 32B variants. These models have been distilled from DeepSeek-R1 and are particularly suitable for demonstrating the effectiveness of reasoning tokens in refining model outputs. To facilitate a rigorous comparison, we also employ these same Qwen2.5 models as baseline (Base) references (without TTC) and in parallel sampling scenarios. Using identical model architectures across conditions ensures a fair and controlled evaluation. 32B models are executed across two NVIDIA A100 GPUs in tensor parallel manner as they on average use over 100GB compared to 80GB memory on our A100 GPUs.
@@
To comprehensively capture the impact of TTC, we evaluate model performance across three representative machine learning task categories, each consisting of four distinct benchmarks. Specifically, we choose mathematical reasoning and code-generation tasks to assess how effectively TTC improves a model’s logical reasoning capabilities. Additionally, common sense tasks are selected to investigate whether extra computation at inference time can meaningfully enhance the model’s retention or representation of factual information. Detailed descriptions of tasks and corresponding datasets are summarized in Table 1.
@@ 4 Results
Figure 2: Normalized energy. The first, second, and third four benchmarks represent math, code, and common-sense, respectively. The dotted horizontal line represents the baseline.
We present our results on Base, MV, and RT. Evaluations are run 10 times with batch size 16, maximum sequence length 32768, and data type Bfloat16 unless stated otherwise. We generate 5 samples when running MV with parallel sampling. Common-sense benchmarks take roughly 10 minutes per evaluation while math and code benchmarks take between 1 to 4 hours depending on the benchmark.
@@ 4.1 Overview
In coding benchmarks, scaling model size offers better accuracy/energy gains at lower model scales than RT. Specifically, transitioning from a Base 1.5B to Base 7B model increases accuracy by 16.8% while reducing energy consumption per query by 40.1%. Conversely, applying RT to the 1.5B model achieves a similar accuracy improvement (17.3%) but increases energy per query by 57.4%. However, when targeting higher accuracy regimes (e.g., Base 32B model versus RT 7B model), RT demonstrates a superior slope of accuracy improvement per unit energy consumed (28.1%/Wh for RT versus 20.7%/Wh for Base). To ensure a fair comparison, RT 7B was chosen to match energy usage with the larger Base model, confirming that RT is more advantageous at higher accuracy demands.
@@
Secondly, the use of MV combined with parallel sampling substantially increases energy consumption without proportionate accuracy improvements. This inefficiency occurs because MV relies on simplistic aggregation, where the probability of correctness, $p$, remains unchanged. Particularly, if $p<0.5$ and we sample 5 generations, using MV with multiple candidates can potentially reduce accuracy due to the binomial distribution of generations. Consequently, a more sophisticated aggregation method is necessary to effectively utilize candidate outputs and justify increased energy consumption.
@@
Figure 3: Power readings during runtime.
@@
On average, MV consumes 2.63$\times$, 2.01$\times$, and 4.69$\times$ more energy than the Base configuration for mathematical, coding, and common-sense benchmarks, respectively. RT demonstrates a more substantial average energy increase, using 3.66$\times$, 10.4$\times$, and 1.08$\times$ more energy than Base for math, code, and common-sense tasks, respectively.
The energy cost of RT, in particular, can escalate dramatically, with a peak increase of 97.48$\times$ for the 7B model on the HumanEval coding benchmark. This extreme rise is largely driven by a substantial growth in the number of generated tokens; RT models produce, on average, 4.4$\times$ more tokens than Base, and in extreme cases up to 46.26$\times$ more tokens.
In contrast, although MV multiplies both input and output tokens by a factor of five, the short output sequences do not fully utilize GPU compute and memory BW during decoding. Consequently, this allows MV to maintain minimal latency overhead despite an increased batch size, resulting in less than a proportional 5$\times$ increase in energy consumption. However, common-sense benchmarks, which generate only a single output token correlating the answer to multiple-choice options, consistently exhibit approximately 5$\times$ energy usage relative to Base.
@@
Our analysis reveals that TTC only marginally affects power usage. Specifically, MV introduces an average increase of approximately 5% in power consumption due to increased batch sizes from parallel sampling. Conversely, RT tends to use slightly less power than Base. This unexpected reduction occurs because RT inference is predominantly decode- and especially attention layer-bound, characterized by extremely long output sequences. The resulting high memory requirements for the KV cache create a bottleneck that saturates GPU memory BW but significantly under-utilizes the GPU compute units.
For instance, RT benchmarks produce an average of 7845 tokens per query, translating to an average KV cache size of about 258MB, even for relatively small (7B) models. Given the memory BW saturation threshold (approximately 100MB data array for A100 GPUs [9307857]), the inference process becomes memory-bound, limiting the effective utilization of GPU computational resources. Consequently, the power usage stabilizes at a level consistent with these memory constraints.

### v2

@@ Abstract
Scaling large language models (LLMs) has driven significant advancements, yet it faces diminishing returns and escalating energy demands. This work explores how test-time compute (TTC) can serve as an energy-efficient complement to conventional scaling strategies by allocating additional computational resources at inference time rather than during training. Specifically, we investigate whether employing TTC can achieve superior accuracy-energy trade-offs compared to simply increasing model size. Our empirical analysis reveals that TTC surpasses traditional model scaling in accuracy/energy efficiency, with notable gains in tasks demanding complex reasoning rather than mere factual recall. Further, we identify a critical interaction between TTC performance and output sequence length, demonstrating that strategically adjusting compute resources at inference time according to query complexity can substantially enhance efficiency. Our findings advocate for TTC as a promising direction, enabling more sustainable, accurate, and adaptable deployment of future language models.
@@ 1 Introduction
Test-time compute (TTC) has emerged as a promising approach to significantly enhance the performance of large language models (LLMs) without the reliance on traditional scaling laws. Conventionally, improving neural language models involved scaling either the number of parameters or expanding the training dataset (2020arXiv200108361K). However, this approach demands substantial computational resources, resulting in considerable energy consumption during both training and inference phases. Furthermore, such scaling is increasingly constrained by the limited availability of sufficiently large and diverse datasets(2022arXiv221104325V), highlighting diminishing returns in terms of resource efficiency.
@@
Recent research addresses this limitation by proposing a new strategy to enhance the performance of LLMs: allocating additional computational resources at inference time, known as TTC. TTC is inspired by human reasoning process, where humans reflect and reason logically to answer challenging questions rather than responding impulsively. Similarly, LLMs should leverage additional computational effort during inference to improve their responses. Prominent examples of TTC methodologies include prompting techniques like Chain-of-Thought (CoT) reasoning  (2022arXiv220311171W; 2023arXiv230510601Y), self-revision methods (2024arXiv240601297K; 2024arXiv240615673L; 2023arXiv230317651M), post-training for reasoning (deepseek; star; rest), and ensemble-like approaches aggregating response from multiple parallel generations to improve the final response (llmonk; getting_50; 2023arXiv230306135I; 2022arXiv220311171W). Despite methodological differences, these strategies collectively emphasize additional computational investment during inference to enhance model accuracy.
@@
The rising computational demands of TTC necessitate careful consideration of associated energy costs. Recent industry reports underscore this concern, with Meta attributing up to 70% of its AI power consumption to inference processes (wu2022sustainable), Google reporting 60% of machine learning energy usage (patterson2022carbon), and AWS indicating inference-related demands accounting for 80-90% of computing resources (barr2019amazon) all without TTC. Moreover, broader data center energy consumption trends indicate a rapid escalation, expected to constitute up to 12% of total U.S. electricity usage by 2028 (shehabi2024united). The environmental impact and financial implications of such intensive energy consumption underline the urgency of evaluating the sustainability and efficiency of TTC approaches.
@@
In this paper, we examine the accuracy-energy trade-offs associated with TTC compared to traditional model scaling. Specifically, we investigate the relationship between accuracy improvements and energy consumption incurred by employing TTC methods versus scaling model sizes across multiple benchmarks. Through comprehensive experiments conducted on Qwen2.5 (qwen2.5) models (1.5B, 7B, 14B, and 32B parameters) using A100 GPUs across mathematical, coding, and common-sense reasoning benchmarks, we derive key insights into the comparative advantages and costs of TTC.
@@
  • TTC demonstrates superior accuracy-energy trade-off improvements over traditional model scaling. Notably, smaller models enhanced with TTC can outperform substantially larger models relying solely on scale.
[…]
@@
  • Complex questions drive models with TTC to allocate greater computational resources during inference, paralleling human cognitive processes in handling challenging tasks.
To the best of our knowledge, this study represents one of the first systematic explorations into the energy implications of TTC for LLM inference. Our analysis provides a foundational understanding of how TTC can serve as a more energy-efficient complement to traditional scaling. We explicitly focus on inference-related impacts, leaving exploration of training-phase considerations as important avenues for future research.
@@ 2 Background
Inference for transformer-based LLMs (llama; qwen2.5; gemini; gpt4; vaswani2017attention) typically involves two sequential stages: prefill and decode. Initially, when an input sequence is provided, it undergoes the prefill stage, where all tokens in the seqeunce are processed simultaneously. This stage computes the first output token and generates a Key-Value cache (KV-cache), which stores contextual information from the input sequence. Following the prefill stage, the model enters the decode stage, wherein it autoregressively generates subsequent tokens one-by-one, leveraging the previous KV-cache.
Each of these stages presents distinct computational bottlenecks. The prefill stage is predominantly compute-bound due to large matrix multiplications resulting from processing all tokens simultaneously. In contrast, the decode stage typically becomes memory bandwidth (BW)-bound because of the frequent accesses to the KV-cache.
@@
Recently, leveraging additional computational resources at inference time—known as TTC—has emerged as an effective strategy for enhancing the reasoning capabilities and overall accuracy of LLMs. We categorize TTC techniques into two categories based on their resource utilization patterns: 1) techniques that increase the input tokens more (llmonk; 2022arXiv220311171W; few_shot) or 2) those that increase output tokens more (deepseek; 2023arXiv230317651M; star; rest). Each category uniquely affects the resource demands of the inference pipeline.
Increasing the number of input tokens primarily impacts the prefill stage as the compute requirement scales quadratically with the length of the input sequence. On the other hand, increasing the number of output tokens predominantly affects the decode stage. Specifically, extending the output length linearly increases both the size of the KV-cache and the number of autoregressive decoding steps, scaling memory BW requirements quadratically.
@@
Early pioneering efforts, such as the study by Strubell et al. (strubell2020energy), began to systematically investigate the energy cost of training transformer models, spurring subsequent studies that quantified energy consumption during model training (patterson2022carbon). With the increasing deployment of LLMs in practical applications, recent studies have expanded this scope to include energy measurements specific to inference workloads (luccioni2024power; fernandez2025energy; wu2025unveiling; patel2024characterizing; stojkovic2024towards).
@@ 3 Methodology
Table 1: List of benchmarks used in this study for Math, Code, and Common Sense tasks
Math Code Common Sense
AIME24 (aime24) HumanEval (humaneval) HellaSwag (hellaswag)
AIME25 (aime25) LiveCodeBench (livecodebench) MMLU (mmlu)
GPQADiamond (gpqa) MBPP (mbpp) GSM8K (gsm8k)
Math500 (math500) CodeForces (codeforces) CommonsenseQA (talmor-etal-2019-commonsenseqa)
In this study, we systematically measure the additional energy consumption associated with TTC methods in the context of LLM inference. To quantify this energy overhead, we base our work on Evalchemy (evalchemy) and utilize SGLang (sglang) for model execution on NVIDIA A100 GPUs (a100), each with a maximum power rating of 500W. We run each benchmark from start to end assuming that all questions are present to be processed to eliminate any scheduling overhead or queuing delay. Power consumption during model inference is monitored using the NVIDIA Management Library (NVML) (nvml). We integrate power usage over the inference period to compute total energy consumption for each model evaluation.
@@
We group TTC approaches into two distinct categories: (1) methods that increase the number of input tokens and (2) methods that increase the number of output tokens. For the first category, we select parallel sampling (llmonk; getting_50; 2023arXiv230306135I; 2022arXiv220311171W) with majority vote (MV) to finalize the answer. For the second category, we adopt the reasoning token (RT) approach (deepseek; star; rest). These models have demonstrated that they can reason through their answers by utilizing reasoning tokens. Unless stated otherwise, we evaluate the energy usage and accuracy of MV and RT against baseline (Base) approach which does not use any TTC methodology.
@@
We base our study on models supported by the DeepSeek-R1 framework (deepseek), selecting four representative models from the Qwen2.5 family (qwen2.5): 1.5B, 7B, 14B, and 32B variants. These models have been distilled from DeepSeek-R1 and are particularly suitable for demonstrating the effectiveness of reasoning tokens in refining model outputs. To facilitate a rigorous comparison, we also employ these same Qwen2.5 models as baseline references (without TTC) and in parallel sampling scenarios. Using identical model architectures across conditions ensures a fair and controlled evaluation. 32B models are executed across two NVIDIA A100 GPUs in tensor parallel manner as they on average use over 100GB compared to 80GB memory on our A100 GPUs.
@@
To comprehensively capture the impact of TTC, we evaluate model performance across three representative machine learning task categories, each consisting of four distinct benchmarks. We choose mathematical reasoning and code-generation tasks to assess how effectively TTC improves a model’s logical reasoning capabilities using mathematical and programming languages following existing studies (gemini; gpt4; qwen2.5). Additionally, common sense tasks are selected to investigate whether extra computation at inference time can meaningfully enhance the model’s retention or representation of factual information. Detailed descriptions of tasks and corresponding datasets are summarized in Table 1.
@@ 4 Results
We present our results on Base, MV, and RT. Evaluations are run 10 times with batch size 16, maximum sequence length 32768, and data type Bfloat16 unless stated otherwise. We generate 5 samples when running MV with parallel sampling. Common-sense benchmarks take roughly 10 minutes per evaluation while math and code benchmarks take between 1 to 4 hours.
@@ 4.1 Overview
Figure 1: Accuracy versus energy per query averaged across four benchmarks in each task. Each dot in the line represent Qwen2.5 1.5B, 7B, 14B, and 32B from left to right, respectively.
@@
In coding benchmarks, scaling model size offers better accuracy/energy gains at lower model scales than RT. Specifically, transitioning from a Base 1.5B to Base 7B model increases accuracy by 16.8% while reducing energy consumption per query by 40.1%. Conversely, applying RT to the 1.5B model achieves a similar accuracy improvement (17.3%) but increases energy per query by 57.4%. However, when targeting higher accuracy regimes (e.g., Base 32B model versus RT 7B model), RT demonstrates a superior slope of accuracy improvement per unit energy consumed (28.1%/Wh for RT versus 20.7%/Wh for Base).
@@
Secondly, the use of MV combined with parallel sampling substantially increases energy consumption without proportionate accuracy improvements. This inefficiency occurs because MV relies on simplistic aggregation, where the probability of correctness, $p$, remains unchanged. Particularly, if $p<0.5$ and we sample 5 generations, using MV with multiple candidates can potentially reduce accuracy due to the binomial distribution of generations. Consequently, a more efficient aggregation method is necessary to better utilize candidate outputs and justify increased energy consumption.
@@
Figure 2: Energy consumption of MV and RT normalized to Base which does not use TTC. Left bars in each color represent MV and right bars represent RT. The first, second, and third sets of four benchmarks represent math, code, and common sense, respectively. The dotted horizontal line represents Base. Note that the y-axis is cut off from 20 to 110.
@@
[…]
Figure 3: Power readings during runtime averaged across four benchmarks in each task.
@@
Our analysis reveals that TTC only marginally affects power usage. Specifically, MV introduces an average increase of approximately 5% in power consumption due to increased batch sizes from parallel sampling. Conversely, RT tends to use slightly less power than Base. This unexpected reduction occurs because RT inference is predominantly decode- and especially attention layer-bound, characterized by extremely long output sequences. The resulting high memory requirements for the KV-cache create a bottleneck that saturates GPU memory BW but significantly under-utilizes the GPU compute units.
For instance, RT benchmarks produce an average of 7845 tokens per query, translating to an average KV-cache size of about 258MB, even for relatively small (7B) models. Given the memory BW saturation threshold (approximately 100MB data array for A100 GPUs (9307857)), the inference process becomes memory-bound, limiting the effective utilization of GPU computational resources. Consequently, the power usage stabilizes at a level consistent with these memory constraints.